OpenStudy (anonymous):

2 + (x + y) = (2 + x) + y is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

OpenStudy (amistre64):

your thoughts?

OpenStudy (anonymous):

Associative property?

OpenStudy (anonymous):

it obviously has to do with addition...

OpenStudy (anonymous):

2 + (x + y) = (2 + x) + y to make this problem look less scary, you can take the ( )'s out 2+x+y=2+x+y it just basically shows you both are equal.. and it wouldnt matter which u added first

Nnesha (nnesha):

distribute property when you have to multiply bracket by outsider number  like this commuative is like 3+ 2 =2+ 3 and symmetric is like 2 + 5 = 7 then 7 = 2+5 and lastly associate is the rule that rfers to grouping like a + (b + c) = (a + b) + c so which one should be

OpenStudy (amistre64):

associative it is

Nnesha (nnesha):

yes you are right @kaylamarie081

OpenStudy (anonymous):

The "Associative Laws" say: * It doesn't matter how you group the numbers when you add. so you are right

OpenStudy (anonymous):

Thank you.. all of you